How would you prepare a 370 mL 10.3% w/v glucose solution in water?
1
Expert's answer
2018-01-25T12:28:07-0500
ρ = m/V – formula of mass concentration. Find mass:
m = ρ × V
m = 0.103 ×370 = 38.11 (g)
Answer
If to dissolve 38.11 grams of glucose in 370 mL of water, 10.3% w/v solution will be prepared.